You decide to invest some money in two different money markets. One account (a) pays 3% interest. The other account pays 2% but offers $20 bonus.write an expression for each account to determine interest and bonus you woul earn in two years. Use y to represent interest and bonus and x to represent the principle invested

Y=x(.03)(2)=6

Answer by stanbon(75887)

You decide to invest some money in two different money markets.
One account (a) pays 3% interest.
I(2) = P*0.03*2 = 0.06P
======
The other account pays 2% but offers $20 bonus.
I(2) = 20 + P(0.02)*2 = 20 + 0.04P
----------------------------------------------
Write an expression for each account to determine interest and bonus you would earn in two years.
-----
Cheers,
Stan H.
-----------